# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@ 3.1. Integral lattice
Following [CIJ18, Proposition 3.2], we use the following equalities: $z^{-\lambda d\lambda}\left\langle\alpha,\beta\right\rangle_{X}=\left\langle z^{-\mu}\alpha,z^{\mu}\beta\right\rangle_{X}$, $\left\langle z^{\rho}\alpha,z^{\rho}\beta\right\rangle_{X}=\left\langle\alpha,\beta\right\rangle_{X}$, and $e^{2\pi\mathbbm{i}\mu}\rho=-\rho e^{2\pi\mathbbm{i}\mu}$. The proof then follows from the equivariant Hirzebruch–Riemann–Roch theorem.
$\chi_{z}(C,D)=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)\operatorname{Td}_{X}$
$=z^{-\lambda\partial\lambda}\frac{1}{(2\pi\mathbbm{i})^{\dim X}}\int_{X}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\int_{X}e^{-\pi\mathbbm{i}\dim(X)/2}e^{\pi\mathbbm{i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C^{\vee}\otimes D)$
$=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\left\langle e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle z^{-\mu}z^{-\rho}e^{\pi\mathbbm{i}\rho}e^{\pi\mathbbm{i}\mu}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle(e^{-\pi\mathbbm{i}}z)^{-\mu}(e^{-\pi\mathbbm{i}}z)^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\right\rangle_{X}$
$=\frac{1}{(2\pi)^{\dim X}}\left\langle\Psi(C)(e^{-\pi\mathbbm{i}}z),\Psi(D)(z)\right\rangle_{X}$
Let us return now to the specific geometry of the relative Grassmann flop of the previous section. Let $X_{+}$ and $X_{-}$ be as in Section 2. Let $\psi_{\pm}$ denote (3.1.1) for these particular spaces.
@@
Define $\mathbb{U}:\mathcal{H}_{X_{-}}\to\mathcal{H}_{X_{+}}$ to be
$\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1},$
where $\mathbb{U}_{H}$ is the map in (2.3.5).
@@
By the Leray–Hirsch theorem, any class in $\hat{H}^{*}_{\mathbb{T}}(X_{-})$ may be decomposed as a sum of classes of the form $\pi_{-}^{*}(\beta)\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C})$ for some $\overline{C}\in K^{0}_{\mathbb{T}}(\overline{X}_{-})$ where the map $\tilde{v}_{-}$ appears in the proof of Corollary 2.10. We may therefore assume without loss of generality that $f(z)=\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))$ and $g(z)=\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))$. Using the same chain of equalities from the previous proof, the projection formula, and (2.3.8), one computes that
$\left\langle\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{-}}$
$=$ $K\cdot\chi_{z}(\overline{C},\overline{D})$
@@
$\left\langle\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C}))(e^{-\pi\mathbbm{i}}z),\mathbb{U}\psi_{-}(\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{D}))(z)\right\rangle_{X_{+}}$
$=$ $K\cdot\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D}))$
@@
$K=\left(z^{-\lambda\partial\lambda}\frac{(2\pi)^{\dim X}}{(2\pi\mathbbm{i})^{\dim B}}\int_{B}\left((-2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{1}\right)\cup\left((2\pi\mathbbm{i})^{\deg_{0}/2}\beta_{2}\operatorname{Td}_{B}\right)\right).$
@@
$\chi_{z}(\overline{C},\overline{D})=\chi_{z}(\overline{\mathbb{FM}}(\overline{C}),\overline{\mathbb{FM}}(\overline{D})).$
The claim follows.
3.2. $I$-functions for Grassmann bundles
In this section, we record $I$-functions for $X_{\pm}$ and the associated abelian quotients. In each case an $I$-function is obtained by modifying the $J$-function of the base $B$.
3.2.1. $r=1$
@@
$X_{+}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\right)$
$X_{-}=$ $\operatorname{tot}\left(\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\right).$
In this case $X_{\pm}$ is a toric bundle over $B$. $I$-functions for split toric bundles were computed by Brown in [Bro14] in terms of the $J$-function for the base $B$. Let $J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}$ be the $J$-function for $B$. Define the $H^{*}_{\mathbb{T}}(X_{\pm})$-valued function $I_{X_{\pm}}(Q,q,z)$ to be
(3.2.1) $I_{X_{+}}(Q,q_{+},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{e\in\mathbb{Z}}q_{+}^{-y/z+e}M_{\beta,-e}$
(3.2.2) $I_{X_{-}}(Q,q_{-},z)=$ $\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{d\in\mathbb{Z}}q_{-}^{y/z+d}M_{\beta,d}$
@@
(3.2.3) $M_{\beta,d}=\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y+hz)}{\prod_{h=-\infty}^{d-\beta\cdot M_{i}}(z_{i}+y+hz)}\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y+hz)}{\prod_{h=-\infty}^{-d+\beta\cdot L_{k}}(-x_{k}-y+hz)}.$
@@
Despite having the same formal expression under the change of variables $q_{+}=q_{-}^{-1}$, we emphasize that $I_{X_{+}}$ and $I_{X_{-}}$ take very different forms due to the fact that they lie in different cohomology rings. In particular, in $I_{X_{+}}$, the expression $M_{\beta,-e}$ vanishes whenever $e<\operatorname{min}_{k}(-\beta\cdot L_{k})$ by (2.2.1). On the other hand in $I_{X_{-}}$ the expression $M_{\beta,d}$ vanishes when $d<\operatorname{min}_{i}(\beta\cdot M_{i})$.
The following is a special case of Brown’s result for toric bundles (see [Bro14]).
@@
3.2.2. The associated abelian quotients
@@
$L_{\beta,\vec{d}}=$ $\prod_{j=1}^{r}\left(\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}.\right)$
Define $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},z)$ to be
(3.2.4) $I_{X_{T,j}}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{\stackrel{{\scriptstyle(e_{1},\ldots,e_{j})\in\mathbb{Z}^{j}}}{{(d_{j+1},\ldots,d_{r})\in\mathbb{Z}^{s-j}}}}\prod_{i=1}^{j}q_{i,+}^{-y_{i}/z+e_{i}}\prod_{i=j+1}^{r}q_{i,-}^{y_{i}/z+d_{i}}L_{\beta,d_{1},\ldots,d_{j},-e_{j+1},\ldots,-e_{r}}.$
Again by [Bro14], we see that $I_{X_{T,j}}(Q,d_{1},\ldots,d_{j},e_{j+1},\ldots,e_{r},-z)$ lies on the Lagrangian cone for $X_{T,j}$.
3.2.3. The non-abelian quotients
The abelian/non-abelian correspondence in Gromov–Witten theory is the principle that one can recover the Gromov–Witten theory of the non-abelian quotient from that of the associated abelian quotient. There is a long history of results in this direction, e.g. [BCFK05, BCFK08, CFKS08, Web23, Web21, CLS22]. The modern formulation of the expected relationship is, roughly, that the (genus zero) Gromov–Witten theory for the non-abelian quotient may be obtained from that of the associated abelian quotient by the following steps:
  1. (1)

### v2

@@ 3.1 Integral lattice
Following [14, Proposition 3.2], we use the following equalities:
$z^{-\lambda{\rm d}\lambda}\left\langle\alpha,\beta\right\rangle_{X}=\left\langle z^{-\mu}\alpha,z^{-\mu}\beta\right\rangle_{X},\left\langle z^{\rho}\alpha,z^{\rho}\beta\right\rangle_{X}=\left\langle\alpha,\beta\right\rangle_{X}\qquad\text{and}\qquad{\rm e}^{2\pi{\rm i}\mu}\rho=-\rho{\rm e}^{2\pi{\rm i}\mu}.$
The proof then follows from the equivariant Hirzebruch–Riemann–Roch theorem,
$\chi_{z}(C,D)=z^{-\lambda\partial\lambda}\frac{1}{(2\pi{\rm i})^{\dim X}}\int_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}\operatorname{Td}_{X}$
$\qquad{}=z^{-\lambda\partial\lambda}\frac{1}{(2\pi{\rm i})^{\dim X}}\int_{X}{\rm e}^{\pi{\rm i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}$
$\qquad{}=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\int_{X}{\rm e}^{-\pi{\rm i}\dim(X)/2}{\rm e}^{\pi{\rm i}\rho}\hat{\Gamma}_{X}^{\vee}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}\bigl{(}C^{\vee}\otimes D\bigr{)}$
$\qquad{}=\frac{z^{-\lambda\partial\lambda}}{(2\pi)^{\dim X}}\bigl{\langle}{\rm e}^{\pi{\rm i}\rho}{\rm e}^{\pi{\rm i}\mu}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}z^{-\mu}z^{-\rho}{\rm e}^{\pi{\rm i}\rho}{\rm e}^{\pi{\rm i}\mu}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)}^{-\mu}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)}^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(C),z^{-\mu}z^{\rho}\hat{\Gamma}_{X}(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(D)\bigr{\rangle}_{X}$
$\qquad{}=\frac{1}{(2\pi)^{\dim X}}\bigl{\langle}\Psi(C)\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\Psi(D)(z)\bigr{\rangle}_{X}.$
Let us return now to the specific geometry of the relative Grassmann flop of the previous section. Let $X_{+}$ and $X_{-}$ be as in Section 2. Let $\psi_{\pm}$ denote (3.1) for these particular spaces.
@@
Define $\mathbb{U}\colon\mathcal{H}_{X_{-}}\to\mathcal{H}_{X_{+}}$ to be $\mathbb{U}=\psi_{+}\circ\mathbb{U}_{H}\circ\psi_{-}^{-1}$, where $\mathbb{U}_{H}$ is the map in (2.8).
@@
By the Leray–Hirsch theorem, any class in $\hat{H}^{*}_{\mathbb{T}}(X_{-})$ may be decomposed as a sum of classes of the form $\pi_{-}^{*}(\beta)\operatorname{ch}^{\mathbb{T}}(\tilde{v}_{-}^{*}\overline{C})$ for some $\overline{C}\in K^{0}_{\mathbb{T}}\bigl{(}\overline{X}_{-}\bigr{)}$, where the map $\tilde{v}_{-}$ appears in the proof of Corollary 2.11.
We may therefore assume without loss of generality that $f(z)=\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}$ and $g(z)=\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}$. Using the same chain of equalities from the previous proof, the projection formula, and (2.11), one computes that
$\bigl{\langle}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}(z)\bigr{\rangle}_{X_{-}}=K\cdot\chi_{z}\bigl{(}\overline{C},\overline{D}\bigr{)}$
@@
$\bigl{\langle}\mathbb{U}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{1})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{C}\bigr{)}\bigr{)}\bigl{(}{\rm e}^{-\pi{\rm i}}z\bigr{)},\mathbb{U}\psi_{-}\bigl{(}\pi_{-}^{*}(\beta_{2})\operatorname{ch}^{\mathbb{T}}\bigl{(}\tilde{v}_{-}^{*}\overline{D}\bigr{)}\bigr{)}(z)\bigr{\rangle}_{X_{+}}$
$\qquad{}=K\cdot\chi_{z}\bigl{(}\overline{\mathbb{FM}}\bigl{(}\overline{C}\bigr{)},\overline{\mathbb{FM}}\bigl{(}\overline{D}\bigr{)}\bigr{)},$
@@
$K=\biggl{(}z^{-\lambda\partial\lambda}\frac{(2\pi)^{\dim X}}{(2\pi{\rm i})^{\dim B}}\int_{B}\bigl{(}(-2\pi{\rm i})^{\deg_{0}/2}\beta_{1}\bigr{)}\cup\bigl{(}(2\pi{\rm i})^{\deg_{0}/2}\beta_{2}\operatorname{Td}_{B}\bigr{)}\biggr{)}.$
@@
$\chi_{z}\bigl{(}\overline{C},\overline{D}\bigr{)}=\chi_{z}\bigl{(}\overline{\mathbb{FM}}\bigl{(}\overline{C}\bigr{)},\overline{\mathbb{FM}}\bigl{(}\overline{D}\bigr{)}\bigr{)}.$
The claim follows. ∎
3.2 $\bm{I}$-functions for Grassmann bundles
In this section, we record $I$-functions for $X_{\pm}$ and the associated abelian quotients. In each case, an $I$-function is obtained by modifying the $J$-function of the base $B$.
3.2.1 $\bm{r=1}$
@@
$X_{+}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(E)}(-1)\otimes p^{*}F^{\vee}\bigr{)},\qquad X_{-}=\operatorname{tot}\bigl{(}\mathcal{O}_{\mathbb{P}(F^{\vee})}(-1)\otimes p^{*}E\bigr{)}.$
In this case $X_{\pm}$ is a toric bundle over $B$. $I$-functions for split toric bundles were computed by Brown in [7] in terms of the $J$-function for the base $B$. Let $J^{B}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}$ be the $J$-function for $B$. Define the $H^{*}_{\mathbb{T}}(X_{\pm})$-valued function $I_{X_{\pm}}(Q,q,z)$ to be
$I_{X_{+}}(Q,q_{+},z)=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{e\in\mathbb{Z}}q_{+}^{-y/z+e}M_{\beta,-e},$
$I_{X_{-}}(Q,q_{-},z)=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{d\in\mathbb{Z}}q_{-}^{y/z+d}M_{\beta,d},$
@@
$M_{\beta,d}=\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y+hz)}{\prod_{h=-\infty}^{d-\beta\cdot M_{i}}(z_{i}+y+hz)}\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y+hz)}{\prod_{h=-\infty}^{-d+\beta\cdot L_{k}}(-x_{k}-y+hz)}.$ (3.3)
@@
Despite having the same formal expression under the change of variables $q_{+}=q_{-}^{-1}$, we emphasize that $I_{X_{+}}$ and $I_{X_{-}}$ take very different forms due to the fact that they lie in different cohomology rings. In particular, in $I_{X_{+}}$, the expression $M_{\beta,-e}$ vanishes whenever $e<\operatorname{min}_{k}(-\beta\cdot L_{k})$ by (2.2). On the other hand, in $I_{X_{-}}$ the expression $M_{\beta,d}$ vanishes when $d<\operatorname{min}_{i}(\beta\cdot M_{i})$.
The following is a special case of Brown’s result for toric bundles (see [7]).
@@
3.2.2 The associated abelian quotients
@@
$L_{\beta,\vec{d}}={}$ $\prod_{j=1}^{r}\Biggl{(}\prod_{i=1}^{n}\frac{\prod_{h=-\infty}^{0}(z_{i}+y_{j}+hz)}{\prod_{h=-\infty}^{d_{j}-\beta\cdot M_{i}}(z_{i}+y_{j}+hz)}\cdot\prod_{k=1}^{n}\frac{\prod_{h=-\infty}^{0}(-x_{k}-y_{j}+hz)}{\prod_{h=-\infty}^{-d_{j}+\beta\cdot L_{k}}(-x_{k}-y_{j}+hz)}\Biggr{)}.$
Define $I_{X_{T,j}}(Q,d_{1},\dots,d_{j},e_{j+1},\dots,e_{r},z)$ to be
$I_{X_{T,j}}=\sum_{\beta\in NE(B)_{\mathbb{Z}}}Q^{\beta}J_{\beta}\sum_{\stackrel{{\scriptstyle(e_{1},\dots,e_{j})\in\mathbb{Z}^{j}}}{{(d_{j+1},\dots,d_{r})\in\mathbb{Z}^{s-j}}}}\prod_{i=1}^{j}q_{i,+}^{-y_{i}/z+e_{i}}\prod_{i=j+1}^{r}q_{i,-}^{y_{i}/z+d_{i}}L_{\beta,d_{1},\dots,d_{j},-e_{j+1},\dots,-e_{r}}.$
Again by [7], we see that $I_{X_{T,j}}(Q,d_{1},{\dots},d_{j},e_{j+1},{\dots},e_{r},-z)$ lies on the Lagrangian cone for $X_{T,j}$.
3.2.3 The non-abelian quotients
The abelian/non-abelian correspondence in Gromov–Witten theory is the principle that one can recover the Gromov–Witten theory of the non-abelian quotient from that of the associated abelian quotient. There is a long history of results in this direction, e.g., [4, 5, 13, 17, 33, 34]. The modern formulation of the expected relationship is, roughly, that the (genus zero) Gromov–Witten theory for the non-abelian quotient may be obtained from that of the associated abelian quotient by the following steps:
  1. 1)